OpenStudy (anonymous):

Triangle ABC is similar to Triangle DEF. AB=2x+4, AC=4x+1, ED=6x squared+14x+4, EF=9x squared+9x+2, DF=12x squared+7x+1. What is BC? Drawing below...

OpenStudy (anonymous):



OpenStudy (yttrium):

Do you know that the corresponding sides of similar triangles are proportional?

OpenStudy (anonymous):

yeah..

OpenStudy (anonymous):

@Yttrium

OpenStudy (yttrium):

Then, you can formulate a formula for BC using ratio and proportion. Can you?

OpenStudy (anonymous):

AB/DE = AC/DF

OpenStudy (anonymous):

@Yttrium

OpenStudy (yttrium):

Use the segment BC because that is what we are going to solve.

OpenStudy (anonymous):

oh yeah, sorry... BC/EF=AC/DF @Yttrium

OpenStudy (yttrium):

Can you plug in the values?

OpenStudy (yttrium):

@ksalimios

OpenStudy (anonymous):

\[BC/9X ^{2} = 4X+1/12X ^{2}+7X\]

OpenStudy (yttrium):

Then BC will be equal to what?

OpenStudy (anonymous):

i don't know

OpenStudy (yttrium):

You know cross-multiplication?

OpenStudy (anonymous):

yeah.. so \[\left( BC \right)\left( 12X ^{2} +7X+1\right)= \left( 4X+1 \right)\left( 9X ^{2}+9X+2 \right)\]

OpenStudy (yttrium):

Yes, but bring down the term beside BC and factor it out

OpenStudy (anonymous):

that's my problem.. the factoring part

OpenStudy (anonymous):

@Yttrium

OpenStudy (yttrium):

Well, this kind of lesson for me is hard to teach. But if you know the foil method of multiplication, it's better. I can teach you quicker.

OpenStudy (anonymous):

can you teach me please?? @Yttrium

OpenStudy (yttrium):

Awww. I need to log out now. :(( Just ping me if you see me online here so I can still teach you about this topic. But anyway, this is your 1st incentive from me. :) \[BC = \frac{ (4x+1)(3x+1)(3x+2) }{ (3x+1)(4x+1) }\]

OpenStudy (anonymous):

thank u @Yttrium .. it means a lot..